# Enhancing the Robustness of LLM-Generated Code: Empirical Study and Framework

### v3

@@ 2.3.2. Analysis Process
For each code snippet, we analyze whether robustness issues exist in the code snippet and what category can it be categorized into. For example, some robustness issues are arise from flaws in Boolean logic expressions affecting control flow, such as missing conditions, incorrect conditions, or inadequate input validation. Using an open coding approach (Strauss and Corbin, 1990), we iteratively categorize these issues, refining the categories as needed. A single snippet may exhibit multiple issues and receive multiple labels. Two authors independently annotate the data, resolving disagreements through discussion and majority consensus.
To ensure accurate assessment, we consider additional contextual information, including reference code, surrounding class context, compilation feedback, and test results. For complex cases where LLM-generated code implements multiple functionalities, each part is individually compared against the reference implementation. The comparison focuses on semantic equivalence rather than surface-level differences; for instance, expressions like “child == null” and “childNode == null” are considered equivalent if they serve the same logical purpose.
@@ 2.3.3. Results
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 2: We identified nine distinct robustness patterns. Statistical analysis of their frequencies revealed that over 90% of these issues are related to missing conditional checks, with Missing Null Checks being the most prevalent.
@@ 2.4.2. Results
A detailed analysis shows that DeepSeekCoder-1.3B exhibits the highest proportion (75.0%) of robustness issues in the first line, while Qwen2.5-Coder-7B has a lower rate (61.3%). Model size also affects the likelihood of first-line robustness issues. As model size increases, the proportion of missing checks decreases. For instance, DeepSeekCoder-1.3B has a 75.0% issue rate, which drops to 70.0% in DeepSeekCoder-6.7B. Similarly, Qwen2.5-Coder-1.5B has a 69.4% rate, reducing to 61.3% in Qwen2.5-Coder-7B. Additionally, DeepSeekCoder models generate a higher proportion of first-line robustness issues compared to Qwen2.5-Coder models, suggesting that differences in training data or model architecture impact robustness of generated code.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 3: Most robustness issues occur in the first line of the generated code, primarily due to missing robustness checks.
@@ 2.5.1. Design
Building on the findings from RQ3, we first identify code lines where missing condition checks lead to robustness issues. For each identified line, we determine the expected insertion point—marked by the first non-whitespace token—and exclude cases where an “if” statement already exists. This filtering yields 250 code snippets for analysis. For each snippet, we capture the model’s logit outputs at the insertion point and retain the top 30 ranked tokens for further examination.
@@ 2.5.2. Results
Figure 6 shows that at positions where an “if” statement is expected, the “if” token consistently ranks within the top five. For Qwen2.5-Coder-1.5B, when the “if” statement is missing, 50.0% of the cases rank “if” second, 31.7% rank it third, and 10% rank it fourth. Similarly, for DeepSeekCoder-6.7B, 43.6% rank “if” second, 29.0% third, and 9.68% fourth.
@@
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 4: The high ranking of the “if” token (usually second or third) at positions lacking conditionals suggests that LLMs inherently recognize essential control structures. Optimizing the sampling strategy could help generate more robust code by adding missing “if” branches.
3. Plug-in Framework for Enhancing Code Robustness
Based on our findings, we propose two plug-in techniques within the RobGen framework to enhance code robustness in LLMs: RobGen-Adj (Section 3.1) and RobGen-Ins (Section 3.2), targeting different phases of the code generation process. Figure 7 provides an overview of the RobGen framework. RobGen-Adj is a plug-in that adjusts token probabilities during decoding to encourage the generation of control structures, while RobGen-Ins is a post-generation plug-in that inserts missing conditionals to enhance robustness. RobGen-Adj and RobGen-Ins can operate independently or be combined as needed. RobGen focuses on addressing robustness issues related to missing conditional checks, which account for more than 90% of occurrences (Finding 2).
@@ 3. Plug-in Framework for Enhancing Code Robustness
Figure 8. RobustCoder Prompt
3.1. RobGen-Adj: Decoding Adjustment
Empirical findings from RQ4 (Section 2.5) show that LLMs inherently recognize essential control structures for robustness, as indicated by the high ranking of the “if” token at key positions. Given that token selection is influenced by logits, we hypothesize that dynamically adjusting these logits can guide the model to generate more robust code, especially through conditional statements like “if” branches.
We propose an intervention mechanism that adjusts the logits of the “if” token, assuming that once the model starts generating an “if” statement, it can complete the logic correctly, enhancing code robustness. During token generation, we determine if logit adjustments are needed based on predefined rules and modify the logits accordingly. RobGen-Adj is implemented using the LogitsProcessor interface provided by Transformer(Wolf et al., 2020).
Rule-based Intervention Identification. Unconditionally modifying the “if” token’s logit value at every generation step may disrupt normal code synthesis. To mitigate excessive intervention, we leverage insights from RQ4 (Finding 4), which indicate that the “if” token typically ranks second or third at positions where a conditional statement is missing. Thus, we selectively adjust the logit value only when the if token appears within the top three ranked logits.
@@ 3.1. RobGen-Adj: Decoding Adjustment
To determine $\Delta$, we analyze RQ4 data by calculating the relative difference between the “if” token’s logit and the highest-ranked token, normalized by their rank difference. The adjustment factor $\Delta$ is set to the 90th-percentile of these observed differences, ensuring it surpasses 90% of the values. The derived $\Delta$ values for different models are 2.10, 2.11, 1.63, and 2.29 for DeepSeekCoder-1.3B, DeepSeekCoder-6.7B, Qwen2.5-Coder-1.5B, and Qwen2.5-Coder-7B, respectively. For models not investigated, $\Delta$ can be set to 1.0.
3.2. RobGen-Ins: Post-Generation Insertion
Based on findings from RQ3 (Section 2.4), we observe that robustness issues are often present in the first line of generated code. To address this, we introduce a post-generation insertion mechanism that ensures missing conditional checks, specifically the inclusion of an “if” statement, are handled. It also follows best practices in defensive programming (McConnell, 2004) by adding necessary checks at the beginning of the method implementation.
Given a code generated by LLMs using a specific prompt, we first verify whether the generated code’s first line includes an appropriate check (i.e., an “if” condition). If this check is absent, we trigger an intervention to insert it. The process proceeds as follows:
Check for Insertion Need. We examine the first line of the generated method implementation. If the line lacks an “if” condition (e.g., a missing robustness check), the system triggers a reconstruction of that line using a “Fill-in-the-Middle” (FIM) technique (Bavarian et al., 2022). If an “if” condition is already present, the process terminates.
Conditional Statement Generation Based on FIM. We construct a FIM prompt consisting of three parts: “prefix,” “middle,” and “suffix.” The model’s task is to fill in the placeholder “middle” section (e.g., ¡FIM_Middle¿) by generating the appropriate “if” statement based on the context provided in the “prefix” and “suffix.” The “prefix” includes the method signature along with relevant task-specific context, while the “suffix” consists of the remaining method body following the first line of the generated code.
To guide the model in generating a valid “if” condition, we pre-insert the “if” keyword at the position where the check should appear. This ensures that the model generates code containing the necessary conditional logic.
Post-Generation Filtering. Once the model generates the “if” statement, we perform a quality check. If multiple conditional blocks (e.g., “if-else”) are generated, we retain only the first complete block, as it is typically the highest quality. Subsequent or incorrect blocks are discarded. The generated “if” statement is evaluated based on predefined rules to ensure its correctness and compatibility with the surrounding code. If it passes the validation, it is merged into the original code. Invalid or incomplete conditions are filtered out, preventing errors or inconsistencies from being introduced into the generated code.
This method ensures that only high-quality, relevant “if” statements are inserted, improving the robustness of the generated code while avoiding noise. Note that different LLMs utilize varying special tokens and concatenation strategies for FIM-based training, their FIM templates differ. We employ model-specific templates as provided in the official documentation.
@@ 4.1.1. Design
We conducted experiments using our framework on the four models from our empirical analysis, as well as on StarCoder2-7B. To evaluate the effectiveness of our method, we include the RobustCoder Prompt as a baseline in our experiment. RobustCoder Prompt (RP) involves modifying the prompt to encourage the model to generate more robust code by explicitly guiding it toward incorporating necessary checks and handling potential exceptions. These robustness requirements are embedded into the model’s code-generation template, as illustrated in Figure 8. In all methods, we default to employing greedy sampling, with the token limit set to 300. For different models, the adjustment factor $\Delta$ used during generation varied. Specifically, for the four models from our empirical analysis, we adopted empirically determined values, whereas for the StarCoder2-7B, we set the step size to 1. After each prefix, we append the string “if” to guide the model in generating an if-branch. Specifically, since FIM-based generation may produce additional code, we truncate the FIM-generated results and retain only the first if-branch.
4.1.2. Robustness Metric Results
The metric calculation results, presented in Table 3, demonstrate that RobGen effectively enhances the robustness of generated code across five models. When applying RobGen-Adj+Ins, the AvgABE of model-generated code exceeds that of the reference code. For instance, using RobGen-Adj on DeepSeekCoder-6.7B increases the AvgABE from 1.45 to 2.99, while applying RobGen-Ins to Qwen2.5-Coder-1.5B raises the AvgABE from 1.97 to 2.67—both surpassing the reference code’s AvgABE of 2.03. Likewise, both RobGen-Adj and RobGen-Ins significantly improve ABES values. For example, after applying RobGen-Ins, the ABES value of StarCoder2-7B increases from 0.41 to 0.48, whereas Qwen2.5-Coder-7B’s ABES rises from 0.43 to 0.50 with RobGen-Adj. Among all methods, RobGen-Adj+Ins achieves the best performance, indicating that RobGen-Ins and RobGen-Adj effectively complement each other.
@@ 4.1.2. Robustness Metric Results
The RRI and Pass@1 distributions in Figure 9 highlight the effectiveness of RobGen. For instance, with RobGen-Adj, the proportion of generated code with RRI ¡ 0 for DeepSeekCoder-1.3B drops from 56.5% to 35.7%. Similarly, for StarCoder2-7B, the proportion of code with RRI ¿ 0 rises from 19.6% to 53.0% using RobGen-Adj+Ins.
In comparison, RobustCoder Prompt (RP) is less effective than RobGen. While RP can enhance the robustness of model-generated code, its impact is inconsistent and, in some cases, even detrimental. For instance, after applying RP to StarCoder2-7B, the ABES value remains unchanged, while for Qwen2.5-Coder-1.5B, the AvgABE value decreases from 1.97 to 1.90. These results suggest that adjusting input prompts alone is insufficient for improving the robustness of model-generated code.
Although RobGen significantly enhances the robustness of generated code, its effect on the Pass@1 metric is not always stable. For example, after applying RobGen-Ins, the Pass@1 score for StarCoder2-7B decreases from 41.7 to 39.1. Similarly, Qwen2.5-Coder-1.5B experiences a decline from 39.13 to 36.96 with RobGen-Adj, and Qwen2.5-Coder-7B drops from 48.7 to 47.0 with RobGen-Adj+Ins. However, RobGen can also improve Pass@1 in certain cases: the Pass@1 of DeepSeekCoder-1.3B increases from 34.4 to 37.4 with RobGen-Adj+Ins, while DeepSeekCoder-6.7B improves from 45.7 to 47.8 with RobGen-Adj. We attribute this variability to differences in how models adapt to each method, leading to inconsistent performance across models.
@@
Table 3. AvgABE and ABSE of Diferent Method: “GS” indicates Greedy Sampling and “PGP” indicates Pre-Generation Prompting
Model GS RP RobGen-Adj RobGen-Ins RobGen-Adj+Ins
@@
DSC-1.3B 1.37 1.69 (+0.32) 2.87 (+1.50) 2.32 (+0.95) 3.32 (+1.95)
DSC-6.7B 1.45 1.72 (+0.27) 2.99 (+1.54) 2.42 (+0.97) 3.34 (+1.89)
QWC-1.5B 1.97 1.90 (-0.07) 5.40 (+3.43) 2.67 (+0.70) 5.65 (+3.68)
QWC-7B 1.76 2.06 (+0.30) 2.93 (+1.17) 2.48 (+0.72) 3.22 (+1.46)
STC-7B 1.83 1.83 (+0.00) 4.27 (+2.44) 2.47 (+0.64) 4.49 (+2.66)
ABSE
DSC-1.3B 0.35 0.38 (+0.03) 0.41 (+0.06) 0.43 (+0.08) 0.44 (+0.09)
DSC-6.7B 0.41 0.45 (+0.04) 0.47 (+0.06) 0.48 (+0.07) 0.49 (+0.09)
QWC-1.5B 0.42 0.43 (+0.01) 0.50 (+0.08) 0.47 (+0.05) 0.50 (+0.08)
QWC-7B 0.43 0.44 (+0.01) 0.50 (+0.07) 0.47 (+0.04) 0.51 (+0.08)
STC-7B 0.41 0.41 (+0.00) 0.48 (+0.07) 0.48 (+0.07) 0.50 (+0.09)
Figure 9. RRI and Pass@1 of Different Methods. RRI ¡ 0: Less robust; RRI ¿ 0: More robust (vs. reference).
4.1.3. Example Demonstration of Method-Generated Code
Figure 10 presents the generated results of DeepSeekCoder-1.3B on task “6367670b1a6d9265ec017a0f” and DeepSeekCoder-6.7B on task “636766fe1a6d9265ec017833” using different methods. In Figure (a), we observe that, compared to the ground truth, the output of Greedy Sampling lacks null pointer and empty array checks. When using RobustCoder Prompt, the model still fails to generate these missing robustness checks. However, when applying RobGen-Adj and RobGen-Ins, the model successfully generates the necessary robustness checks, as highlighted in the red box. This demonstrates that RobGen-Adj and RobGen-Ins are more effective than RP in enhancing code robustness.
In Figure (b), we note that the code generated by Greedy Sampling deviates significantly from the Ground Truth. With RobustCoder Prompt, the generated code aligns more closely with the Ground Truth, as RobustCoder Prompt prioritizes avoiding unnecessary try-catch blocks for exception handling. When applying RobGen-Adj, the model-generated code not only fully aligns with the Ground Truth but also includes an additional robustness check to verify file existence before executing subsequent logic. However, with RobGen-Ins, the model introduces a file existence check but fails to fully align with the Ground Truth. This suggests that, compared to RobGen-Ins, RobGen-Adj offers greater flexibility, as it not only aids in generating robustness checks but also guides the model toward producing more accurate and reliable code.

### v4

@@ 2.3.2. Analysis Process
For each code snippet, we analyze whether robustness issues exist in the code snippet and what category can it be categorized into. For example, some robustness issues are arise from flaws in Boolean logic expressions affecting control flow, such as missing conditions, incorrect conditions, or inadequate input validation. Using an open coding approach (Strauss and Corbin, 1990), we iteratively categorize these issues, refining the categories as needed. A single snippet may exhibit multiple issues and receive multiple labels. Two authors independently annotate the data, resolving disagreements through discussion and majority consensus.The computed kappa agreement score is 0.71.
To ensure accurate assessment, we consider additional contextual information, including reference code, surrounding class context, compilation feedback, and evaluation outputs of the LLM evaluator. For complex cases where LLM-generated code implements multiple functionalities, each part is individually compared against the reference implementation. The comparison focuses on semantic equivalence rather than surface-level differences; for instance, expressions like “child == null” and “childNode == null” are considered equivalent if they serve the same logical purpose.
@@ 2.3.3. Results
Finding 2: We identified nine distinct robustness patterns. Statistical analysis of their frequencies revealed that over 90% of these issues are related to missing conditional checks, with Missing Null Checks being the most prevalent.
@@ 2.4.2. Results
A detailed analysis shows that DeepSeekCoder-6.7B exhibits the highest proportion (79.6%) of robustness issues in the first line, while Qwen2.5-Coder-7B has a lower rate (75.6%). Model size also affects the likelihood of first-line robustness issues. As model size increases, the proportion of missing checks increases. For instance, DeepSeekCoder-1.3B has a 73.0% issue rate, which grows to 79.6% in DeepSeekCoder-6.7B. Similarly, Qwen2.5-Coder-1.5B has a 65.4% rate, rising to 75.6% in Qwen2.5-Coder-7B. Additionally, DeepSeekCoder models generate a higher proportion of first-line robustness issues compared to Qwen2.5-Coder models, suggesting that differences in training data or model architecture impact robustness of generated code.
Finding 3: Most robustness issues occur in the first line of the generated code, primarily due to missing robustness checks.
@@ 2.5.1. Design
Building on the findings from RQ3, we first identify code lines where missing condition checks lead to robustness issues. For each identified line, we determine the expected insertion point—marked by the first non-whitespace token—and exclude cases where an “if” statement already exists. This filtering yields 194 code snippets for analysis. For each snippet, we capture the model’s logit outputs at the insertion point and retain the top 30 ranked tokens for further examination.
@@ 2.5.2. Results
Figure 7 shows that at positions where an “if” statement is expected, the “if” token consistently ranks within the top five. For Qwen2.5-Coder-7B, when the “if” statement is missing, 42.9% of the cases rank “if” second, 20.0% rank it third, and 11.4% rank it fourth. Similarly, for DeepSeekCoder-6.7B, 50.0% rank “if” second, 35.4% third, and 6.3% fourth.
@@
Finding 4: The high ranking of the “if” token (usually second or third) at positions lacking conditionals suggests that LLMs inherently recognize essential control structures. Optimizing the sampling strategy could help generate more robust code by adding missing “if” branches.
3. Plug-and-Play Framework for Enhancing Code Robustness
Building on our empirical findings, we present RobGen, a plug-and-play framework that can be seamlessly integrated into the decoding process of LLM-based code generation to improve the robustness of the generated code. The underlying model still performs standard next-token prediction in an auto-regressive manner. RobGen intervenes at the beginning of each line—specifically, at the first non-empty token—where a checker examines the previously generated context and decides whether logit adjustment is required. If adjustment is necessary, the strategy is applied before decoding continues. This procedure repeats until an ¡EOS¿ token is generated or the maximum token limit is reached. Figure 8 illustrates the overall architecture of RobGen, which consists of two key components: (i) a checker that determines when intervention is needed (Section 3.1), and (ii) a logit adjustment module that specifies how the intervention is applied (Section 3.2). Together, these components address common robustness issues, particularly missing conditional checks, which account for over 90% of the observed failures (Finding 2).
3.1. Line-Level Intervention Checker
Uncontrolled logit adjustment at every decoding step may introduce redundant or nested conditional branches. To address this, we design a line-level intervention checker that determines whether the upcoming line should begin with an “if” statement. We adopt a line-level granularity because intervening at every token is both unnecessary and can disrupt normal code generation. By focusing on the first non-whitespace token of each line, the checker guides the model to insert essential control structures, thereby improving the robustness of the generated code without excessive interference. Before generating each new line, the checker receives the previously generated code, including the method signature, and predicts whether a conditional statement is required. If the prediction is positive, the logit adjustment strategy described in Section 3.2 is applied to the first non-whitespace token; otherwise, decoding proceeds without intervention.
To train the checker effectively, we constructed a dataset to teach it when a new line should initiate a conditional statement.
Training Data. We selected three Java repositories from GitHub with over 3,000 stars (Li et al., 2024b)to ensure that the collected methods are of high quality and representative of real-world coding practices. All methods were extracted and segmented into code blocks paired with their method signatures. To provide supervision for the checker, we labeled code blocks as follows: a block is considered a positive sample if the next line in the original code begins with an “if” statement, and a negative sample if it does not. In this context, every code block is required to start with a method signature. This labeling strategy allows the checker to learn when a conditional statement is required at the beginning of a new line. Given that positive examples are naturally less frequent and we aim to avoid unnecessary interventions during decoding, we intentionally include more negative samples. To reduce bias and improve training stability, we applied an imbalanced sampling strategy, resulting in 12,000 samples (4,000 positive, 8,000 negative).
Training Setup. For efficient and effective training, we used Qwen3.0-0.6B (Yang et al., 2025b) as the base model and fine-tuned it with LoRA (Hu et al., 2021) using supervised fine-tuning (SFT) at a learning rate of 2e-5 for five epochs. The model was evaluated on a validation set to monitor generalization performance. The checkpoint with the highest validation accuracy was retained as the final LoRA weights. This setup balances computational efficiency with model adaptation capacity, ensuring the checker can accurately predict when conditional statements should be inserted while keeping training overhead manageable. Other hyperparameters followed the configuration in (Rob, 2025).
@@ 3.1. Line-Level Intervention Checker
3.2. Token-Level Conditional Logit Adjustment
Motivated by empirical findings from RQ4 (Section 2.5), we observe that LLMs generally rank the “if” token highly at positions where a conditional statement is required. Since token selection depends on logits, we propose dynamically adjusting the logits of the “if” token to guide the model toward generating missing conditional statements, thereby enhancing code robustness.
We implement a selective logit adjustment mechanism for the “if” token. By boosting the logit of “if” at strategic positions, the model is guided to generate necessary conditional statements without interfering with overall code synthesis. The mechanism is applied during decoding using the LogitsProcessor interface from Transformers (Wolf et al., 2020).
Selective Intervention Rule. Unconditional logit adjustment at every decoding step may disrupt normal code generation. Based on insights from RQ4 (Finding 4), the “if” token typically ranks second or third at positions where a conditional statement is missing. Therefore, we apply logit adjustment only when the “if” token is among the top three ranked tokens, ensuring targeted and effective intervention.
@@ 3.2. Token-Level Conditional Logit Adjustment
To determine $\Delta$, we analyze RQ4 data by calculating the relative difference between the “if” token’s logit and the highest-ranked token, normalized by their rank difference. The adjustment factor $\Delta$ is set to the 90th-percentile of these observed differences, ensuring it surpasses 90% of the values. The derived $\Delta$ values for different models are 1.60, 1.70, 1.29, and 1.78 for DeepSeekCoder-1.3B, DeepSeekCoder-6.7B, Qwen2.5-Coder-1.5B, and Qwen2.5-Coder-7B, respectively. For models not investigated, $\Delta$ can be set to 1.0.
@@ 4.1.1. Design
We conducted experiments using our framework on the four models from our empirical analysis, as well as on StarCoder2-7B. To assess the effectiveness of our approach, we consider the following baseline methods.
  • Greedy Sampling: The model’s default code generation approach.
  • Robust Coder Prompt: Robust Coder Prompt (RP) involves modifying the prompt to encourage the model to generate more robust code by explicitly guiding it toward incorporating necessary checks and handling potential exceptions. These robustness requirements are embedded into the model’s code-generation template, as detailed in (Rob, 2025).
  • Post Generation Insertion: Based on RQ3 (Section 2.4), We propose a post-generation insertion (PGI) mechanism to address robustness issues in the first line of generated code. Missing “if” statements are reconstructed using a Fill-in-the-Middle (FIM) prompt (Bavarian et al., 2022), guiding the model to generate the necessary conditional logic.
  • RobGen without Checker :When the Checker is not used, the model performs token-level conditional logit adjustment at every token generation step.
In all methods, we default to employing greedy sampling, with the token limit set to 1024. All metrics are computed over the full dataset; this differs from the configuration used in previous empirical research.
For different models, the adjustment factor $\Delta$ used during generation varied. Specifically, for the four models from our empirical analysis, we adopted empirically determined values, whereas for the StarCoder2-7B, we set the step size to 1.
4.1.2. Results
The metric calculation results, presented in Table 3, demonstrate that RobGen effectively enhances the robustness of generated code across five models. While RobGen achieves the highest average Pass@1(43.57) and average AvgABE(2.42) slightly lower than RobGen without Checker, it demonstrates superior stability, maintaining the model’s Pass@1 with minimal fluctuation. When applying RobGen, the AvgABE of model-generated code exceeds that of the reference code. For instance, using RobGen on DeepSeekCoder-6.7B increases the AvgABE from 1.47 to 2.15, while applying RobGen to Qwen2.5-Coder-1.5B raises the AvgABE from 1.38 to 2.42—both surpassing the reference code’s AvgABE of 2.03. Notably, StarCoder2-7B achieves a very high AvgABE score (7.76) under the RobGen approach without the checker, indicating that StarCoder2-7B generates numerous redundant “if” conditional expressions, which in turn negatively impact the functional correctness of the code. In contrast, when the checker is employed in the RobGen approach, this abnormal behavior of StarCoder2-7B is mitigated—or even eliminated—demonstrating that the checker imposes appropriate constraints on the logit adjustments.
@@ 4.1.2. Results
The Robustness Evaluation Results and Pass@1 distributions in Figure 9 highlight the effectiveness of RobGen. For instance, with RobGen, the proportion of generated code underperform Human for DeepSeekCoder-1.3B drops from 42.2% to 28.7%. Similarly, for StarCoder2-7B, the proportion of generated code outperform Human code rises from 19.1% to 38.3% using RobGen.
In comparison, RobustCoder Prompt (RP) is less effective than RobGen. Although RP can improve code robustness, its effects are inconsistent and sometimes detrimental—for example, Qwen2.5-Coder-1.5B sees AvgABE drop from 1.38 to 1.05 and the proportion of code outperforming human-written code fall from 19.1% to 12.2%, indicating that prompt adjustment alone is insufficient to enhance robustness. While Post-Generation Insertion (PGI) improves code robustness, its impact on Pass@1 is inconsistent (e.g., StarCoder2-7B: 42.2 → 38.7; Qwen2.5-Coder-1.5B: 48.7 → 45.7). In contrast, RobGen enhances robustness without compromising Pass@1 and can even improve it (e.g., DeepSeekCoder-1.3B: 34.4 → 38.7; DeepSeekCoder-6.7B: 45.7 → 47.8). Without the checker, RobGen may reduce Pass@1 (e.g., StarCoder2-7B: 42.2 → 40.9), underscoring the checker’s effectiveness.
@@
Table 3. AvgABE and Pass@1 of Diferent Method: “GS” indicates Greedy Sampling, “RP” indicates RobustCoder Prompt and “PGI” indicates Post-Generation Insertion
Model GS RP PGI RobGen w/o Checker RobGen
@@
DSC-1.3B 1.47 1.83 2.43 2.5 2.19
DSC-6.7B 1.47 1.85 2.43 2.15 2.33
QWC-1.5B 1.38 1.05 2.26 2.42 2.19
QWC-7B 1.79 2.08 2.53 2.24 2.37
STC-7B 2.13 1.85 3.07 7.76 3.03
Avg 1.65 1.73 2.06 3.41 2.42
Pass@1
DSC-1.3B 34.35 33.48 36.52 36.52 38.70
DSC-6.7B 45.65 43.91 45.65 48.26 47.83
QWC-1.5B 39.57 32.61 35.65 39.57 39.13
QWC-7B 48.70 47.39 45.65 49.13 49.13
STC-7B 42.17 46.09 38.70 40.87 43.04
Avg 42.09 40.70 40.43 42.87 43.57
Figure 9. Robustness Evaluation Results and Pass@1 of Different Methods.
Figure 10 presents the generated results of DeepSeekCoder-1.3B and StarCoder2-7B using different methods. In Figure (a), we observe that, compared to the ground truth, the output of Greedy Sampling lacks null pointer and empty array checks. When using RobustCoder Prompt, the model still fails to generate these missing robustness checks. However, when applying PGI and RobGen, the model successfully generates the necessary robustness checks, as highlighted in the red box. This demonstrates that PGI and RobGen are more effective than RP in enhancing code robustness.
In Figure (b), We note that when the code is generated by Greedy Sampling, RP, and PGI, it corresponds to the correct version. However, with RobGen w/o Checker, an unnecessary check of “ current instanceof InputStreamIterator” is introduced, leading to redundant and incorrect operational logic. In contrast, when using RobGen with Checker, the Checker successfully prevents the introduction of such incorrect logic, demonstrating its effective constraining role in logit adjustment.
Finding 5: RobGen attains the highest average Pass@1 and offers improved stability, boosting code generation robustness while maintaining or even enhancing the model’s Pass@1, with the Checker serving as a key enforcing mechanism.